# scale cd/.style=every label/.append style=scale=#1, cells=nodes=scale=#1

### v1

@@ 6. Extended tensor products and tensor products of $p$-permutation bimodules
[…]
    ${}^{(g,h)}(\phi,(Q,f))=(c_{g}\phi c_{h}^{-1},{}^{h}(Q,f)),$
@@ 7. Brauer pairs for $p$-permutation modules and $p$-permutation equivalences
Let $A$ and $B$ be sums of blocks of $kG$ and $kH$, respectively. Boltje and Perepelitsky in [4] introduced the notion of a Brauer pair of a $p$-permutation equivalence $\gamma$ (to be precise, they initially defined the notion for an element $\gamma\in T^{\Delta}(A,B)$ for which $\gamma\cdot_{B}\gamma^{*}=[A]$).
Because a $p$-permutation equivalence can be written as a difference of bimodules whose indecomposable direct summands have twisted diagonal vertices, applying the Brauer construction at a subgroup $X\leq G\times H$ to a $p$-permutation equivalence will only be nonzero when $X$ is twisted diagonal. However, this is not a sufficient criteria for the result to be nonzero, and this motivates the definitions given in this section. We review the definitions and results in [4].
@@
[4, Definition 5.1] Let $M\in{}_{kG}\textbf{triv}$. We call a Brauer pair $(P,e)$ of $kG$ an $M$-Brauer pair if $M(P,e):=e\cdot M(P)\neq 0$, where $M$ has $k[N_{G}(P,e)]e$-module structure. We denote the set of Brauer pairs of $M$ by $\mathcal{BP}(M)$.
@@
[4, Proposition 5.3, 5.4] Let $M,N\in{}_{kG}\mathbf{triv}$.
@@
    Let $M$ and $N$ be indecomposable $p$-permutation $kG$-modules, suppose that $(P,e)\in\mathcal{BP}(kG)$ is both a maximal $M$-Brauer pair and a maximal $N$-Brauer pair, and set $I:=N_{G}(P,e)$. Then $M(P,e)$ and $N(P,e)$ are indecomposable $p$-permutation $kIe$-modules. Moreover, $M\cong N$ if and only if $M(P,e)\cong N(P,e)$ as $kIe$-modules.
@@
[4, Definition 9.6] Let $\omega\in T(kG)$. A Brauer pair $(P,e)$ of $kG$ is called an $\omega$-Brauer pair if $\omega(P,e)=e\omega(P)\neq 0$ in $T(k[N_{G}(P,e)])$. The set of $\omega$-Brauer pairs is denoted $\mathcal{BP}(\omega)$.
Let $\gamma\in T^{\Delta}(A,B)$ be a $p$-permutation equivalence for $A$ and $B$ and let $\Delta(P,\phi,Q)\subset G\times H$ be a twisted diagonal subgroup. Via restriction, we can also view $\gamma(\Delta(P,\phi,Q))\in T(k[N_{G\times H}(\Delta(P,\phi,Q))])$ as an element of the corresponding Grothendieck group of $C_{G}(P)\times C_{H}(Q)=C_{G\times H}(\Delta(P,\phi,Q))$.
@@
The following proposition gives convenient reformulations of being a $\gamma$-Brauer pair.
@@
[4, Proposition 10.8] Let $(\Delta(P,\phi,Q),e\otimes f^{*})\in\mathcal{BP}^{\Delta}(A,B)$ and set $I:=N_{G}(P,e)$ and $J:=N_{H}(Q,f)$. The following are equivalent:
@@
[4, Corollary 10.9] Let $(P,e)$ be an $A$-Brauer pair and define $\Lambda$ as in Theorem 6.7 and let $\Lambda_{B}\subseteq\Lambda$ be the set of pairs $(\psi,(Q,f))$ such that $(Q,f)$ is a $B$-Brauer pair. There exists a unique $H$-orbit of pairs $(\psi,(Q,f))\in\Lambda_{B}$ with the property that $(\Delta(P,\psi,Q),e\otimes f^{*})$ is a $\gamma$-Brauer pair.
@@
[4, Theorem 10.10] Suppose $\gamma$ is a $p$-permutation equivalence for $A$ and $B$. Let $\mathcal{I}$ denote the set of primitive idempotents in $Z(A)$ and let $\mathcal{J}$ denote the set of primitive idempotents in $Z(B)$.
@@
For claims (a) and (b), see [4, Theorem 10.10]. For (c), let $\lambda:=(\Delta(P,\phi,Q),b\otimes c^{*})$ be a Brauer pair for $k[G\times H]$. Then
@@
[4, Proposition 11.1] Let $(\Delta(P,\phi,Q),e\otimes f^{*})$ be a $\gamma$-Brauer pair and set $I:=N_{G}(P,e)$ and $J:=N_{H}(Q,f)$. For every $g\in I$ there exists a unique element $hC_{H}(Q)\in J/C_{H}(Q)$ such that $c_{g}\circ\phi=\phi\circ c_{h}:Q\to P$. Similarly for every $h\in J$, there exists a unique element $gC_{G}(P)\in I/C_{G}(P)$ such that $c_{g}\circ\phi=\phi\circ c_{h}$. These associations define mutually inverse group isomorphisms between $I/C_{G}(P)$ and $J/C_{H}(Q)$. The isomorphism $J/C_{H}(Q)\cong I/C_{G}(P)$ restricts to an isomorphism $QC_{H}(Q)/Q\cong PC_{G}(P)/P,hC_{H}(Q)\mapsto\phi(h)C_{G}(P)$, for $h\in Q$. The group $Y:=N_{G\times H}(\Delta(P,\phi,Q),e\otimes f^{*})$ satisfies $p_{1}(Y)=I,p_{2}(Y)=J,k_{1}(Y)=C_{G}(P),k_{2}(Y)=C_{H}(Q)$.
@@
[4, Theorem 11.4] Let $(\Delta(P,\phi,Q),e\otimes f^{*})\in\mathcal{BP}(\gamma)$ be a $\gamma$-Brauer pair and set $I:=N_{G}(P,e)$, $J:=N_{H}(Q,f)$. Suppose $C_{G}(P)\leq S\leq I$ and $C_{H}(Q)\leq T\leq J$ are intermediate groups related via the isomorphism in the previous proposition, and set $Y:=N_{S\times H}(\Delta(P,\phi,Q),e\otimes f^{*})=N_{S\times T}(\Delta(P,\phi,Q))$. Then the element $e\gamma(P,\phi,Q)f\in T(kY(e\otimes f^{*}))$ after restriction to $Y$ satisfies
@@
[4, Proposition 11.5] Let $(\Delta(D,\phi,E),e_{D}\otimes f^{*}_{E})$ be a maximal $\gamma$-Brauer pair. In particular, $D$ is a defect group of $A$ and $E$ is a defect group of $B$. For any subgroup $P\leq D$, let $(P,e_{P})$ denote the unique $A$-Brauer pair with $(P,e_{P})\leq(D,e_{D})$ and for any subgroup $Q\leq E$ let $(Q,f_{Q})$ denote the unique $B$-Brauer pair with $(Q,f_{Q})\leq(E,f_{E})$. Let $Q\leq E$ and set $P:=\phi(Q)$. Then $(\Delta(P,\phi,Q),e_{P}\otimes f^{*}_{Q})$ is a $\gamma$-Brauer pair.
@@
8. Brauer pairs for splendid Rickard complexes
Each splendid Rickard complex $C$ induces a $p$-permutation equivalence $\gamma:=\Lambda(C)$, and therefore has an associated set of $\gamma$-Brauer pairs. As with $p$-permutation equivalences, in general applying the Brauer construction at an arbitrary subgroup to $C$ will not yield a local splendid Rickard complex. We may define a Brauer pair for a splendid Rickard complex in the same way as we have for $p$-permutation equivalences, treating contractible complexes as zero objects. In fact, the Lefschetz invariant of a contractible complex is zero in the corresponding Grothendieck group, but the converse of the statement rarely holds.
We will show that applying the Brauer construction at the Brauer pairs for the corresponding $p$-permutation equivalence $\gamma$ do indeed yield local splendid equivalences. Moreover, we demonstrate that applying the Brauer construction at any non-Brauer pair results in a contractible complex. The main result of this paper is the corollary of the two above results, which is that the Brauer pairs for a splendid Rickard complex and its corresponding $p$-permutation equivalence are the same. For this section, suppose $A,B$ are direct summands of group algebras $kG$ and $kH$ respectively. Let $C$ be a splendid Rickard complex inducing a splendid equivalence between $A$ and $B$, and set $\gamma:=\Lambda(C)$. Given a Brauer pair $\omega=(\Delta(P,\phi,Q),e\otimes f^{*})$, denote by $\omega^{\circ}$ the “opposite” Brauer pair $(\Delta(Q,\phi^{-1},P),f\otimes e^{*}).$
@@ 8. Brauer pairs for splendid Rickard complexes
Let $C$ be a splendid Rickard complex. The goal of this section is to determine the set $\mathcal{B}\mathcal{P}(C)$ of $C$-Brauer pairs. Notice that the Lefschetz invariant of $C(\omega)$ is $\gamma(\omega)$. Since the Lefschetz invariant of a contractible complex is zero, it is immediate that $\mathcal{BP}(\gamma)\subseteq\mathcal{BP}(C)$. However, a noncontractible complex can have Lefschetz invariant equal to zero, so it is not clear whether or not these sets are equal.
@@
First, we wish to reduce to the case that $A$ and $B$ are blocks. In order to do this, we will establish an analogue of Proposition 7.7 for splendid Rickard complexes. The main tool is the following result.
@@
Since $C$ is a splendid Rickard equivalence, we have a homotopy equivalence $C\otimes_{B}C^{*}\simeq A$ of complexes of $(A,A)$-bimodules. We apply the functor $-(\Delta(P),e\otimes e^{*})$ to this equation. Using Proposition 9.3, Theorem 6.7(d), and Proposition 3.4, we obtain an isomorphism
@@
Suppose $A$ and $B$ are sums of blocks and that $C$ is a splendid Rickard complex for $A$ and $B$. Let $f:\mathcal{I}\to\mathcal{J}$ be the bijection determined by $\gamma:=\Lambda(C)$ from Proposition 7.7.
@@
    We have $C=\bigoplus_{e\in\mathcal{I}}eCf_{e}$ in $K^{b}({}_{A}\mathbf{mod}_{B})$. Moreover, for every $e\in\mathcal{I}$, $eCf_{e}$ is a splendid Rickard complex for $kGe$ and $kGf_{e}$.
@@
The Lefchetz invariant of $eCf_{e}$ is $e\gamma f_{e}$, which is nonzero by Proposition 7.7(a). Therefore, $eCf_{e}\not\simeq 0$. By applying Lemma 8.3 to the splendid Rickard complex $C$ and the $A$-Brauer pair $(1,e)$, we obtain that there exists a unique $f\in\mathcal{J}$ such that $eCf\not\simeq 0$. By uniqueness, $f=f_{e}$. This proves (a).
@@
Therefore, it suffices to consider only splendid Rickard complexes for block algebras. From here, assume $A$ and $B$ are block algebras. We next prove a chain complex-theoretic version of Theorem 7.10.
@@
Suppose that $A$ and $B$ are blocks and $C$ is a splendid Rickard complex for $A$ and $B$, and set $\gamma:=\Lambda(C)$. Set $\lambda:=(\Delta(P,\phi,Q),e\otimes f^{*})$ be a $\gamma$-Brauer pair and set $I:=N_{G}(P,e),J:=N_{H}(Q,f)$. Suppose $C_{G}(P)\leq S\leq I$ and $C_{H}(Q)\leq T\leq J$ are intermediate groups related via the isomorphism $I/C_{G}(P)\to J/C_{H}(Q)$ defined in Proposition 7.9, and set $Y:=N_{S\times H}(\Delta(P,\phi,Q),e\otimes f^{*})$. Then the chain complex $C(\lambda):=eC(\Delta(P,\phi,Q))f\in\mathrm{Ch}^{b}({}_{kY(e\otimes f^{*})}\mathbf{triv})$ satisfies
@@
In particular, $C(\lambda)$ is a splendid Rickard complex for $k[C_{G}(P)]e$ and $k[C_{H}(Q)]f$, and the chain complex
@@
is a splendid Rickard complex for $kSe$ and $kTf$.
@@
of chain complexes of $k[N_{S\times S}(\Delta(P))](e\otimes e^{*})$-modules. Notice that when $\tau=(\phi,(Q,f))$, we have $X(\tau)=Y$, $C(\tau)=C(\lambda)$, and, by Proposition 7.9, $I(\tau):=N_{(S,\phi,J)}=S$. It follows that from the latter that
@@
Thus, $C(\lambda)\otimes_{H}^{Y,Y^{\circ}}C(\lambda)^{*}$ is a noncontractible direct summand of the complex on the right hand side. On the other hand, $k[C_{G}(P)]e$ is an indecomposable $k[N_{S\times S}(\Delta(P))]$-bimodule, hence indecomposable and noncontractible as a chain complex. By the Krull-Schmidt theorem, precisely one of the terms in the direct sum must be noncontractible. We conclude that $k[C_{G}(P)]e\simeq C(\lambda)\otimes_{H}^{Y,Y^{\circ}}C(\lambda)^{*}$. Therefore,
@@
The second homotopy equivalence follows similarly, and therefore $\operatorname{Ind}^{S\times T}_{Y}(C(\lambda))$ is a splendid Rickard complex for $kSe$ and $kTf$. ∎
We next show that in addition, if $\omega:=(\Delta(P,\phi,Q),e\otimes f^{*})$ is not a $\gamma$-Brauer pair, then $C(\omega)\simeq 0$ as a complex of $k[N_{G\times H}(\omega)]$-modules.
@@
    $\operatorname{Res}^{G}_{H}\operatorname{Ind}^{G}_{H}C\cong\bigoplus_{g\in[H\backslash G/H]}\operatorname{Ind}^{H}_{H\cap{}^{g}H}\operatorname{Res}^{H}_{H\cap{}^{g}H}C.$
    Suppose $\operatorname{Ind}^{G}_{H}C\simeq 0$. Then the left-hand side is contractible since restriction preserves contractibility. All summands on the right-hand side are contractible as well, since direct summands of contractible chain complexes are again contractible. Since there is a coset representative in $[H\backslash G/H]$ belonging to $H$, $C$ is isomorphic to a direct summand of $\operatorname{Res}^{G}_{H}\operatorname{Ind}^{G}_{H}C$, and thus is contractible. The reverse direction is clear since $\operatorname{Ind}_{H}^{G}$ is an additive functor.
@@
Suppose $C$ is a splendid Rickard complex for $A$ and $B$, with $\gamma=\Lambda(C)$. If $\omega:=(\Delta(R,\psi,Q),d\otimes f^{*})\in\mathcal{BP}^{\Delta}(A,B)$ is not a $\gamma$-Brauer pair, then $C(\omega):=dC(\Delta(R,\psi,Q))f$ is contractible as a chain complex of $k[N_{G\times H}(\Delta(R,\psi,Q),d\otimes f^{*})](d\otimes f^{*})$-modules. In particular, $\omega$ is not a $C$-Brauer pair.
@@
There exists an $A$-Brauer pair $(P,e)$ and an isomorphism $\phi:Q\to P$ such that $\lambda:=(\Delta(P,\phi,Q),e\otimes f^{*})$ is a $\gamma$-Brauer pair. Set $I=N_{G}(P,e),\ J=N_{H}(Q,f)$, $K=N_{G}(R,d)$, $X=N_{G\times H}(\lambda)$, and $Y=N_{G\times H}(\omega)^{\circ}$. By Theorem 8.5, $\operatorname{Ind}_{X}^{I\times J}(C(\lambda))$ is a splendid Rickard complex for $kIe$ and $kJf$. In order to prove the claim, it suffices to show that
@@
$\operatorname{Ind}_{X}^{I\times J}(C(\lambda))\mathop{\otimes}\limits_{kJ}\operatorname{Ind}_{Y}^{J\times K}(C(\omega)^{*})\cong\bigoplus_{t\in[p_{2}(X)\backslash J/p_{1}(Y)]}\operatorname{Ind}_{X\ast{}^{(t,1)}Y}^{I\times K}\left(C(\lambda)\mathop{\otimes}\limits_{kJ}^{X,{}^{(t,1)}Y}{}^{(t,1)}(C(\omega)^{*})\right).$

### v2

@@ 6. Extended tensor products and tensor products of $p$-permutation bimodules
    Assume that $G=K$, $(P,e)=(R,d)$, $S=T$, $\sigma=\text{id}_{P}$. Let $\Lambda:=\Lambda_{H}(P)$ be the set of pairs $(\phi,(Q,f))$ where $(Q,f)$ is a Brauer pair of $kH$ and $\phi:Q\to P$ is an isomorphism. The group $S\times H$ acts on $\Lambda$ via
    $),{}^{(g,h)}(\phi,(Q,f))=(c_{g}\phi c_{h}^{-1},{}^{h}(Q,f$)),
@@ 7. Brauer pairs for $p$-permutation modules and $p$-permutation equivalences
Let $A$ and $B$ be sums of blocks of $kG$ and $kH$, respectively. Boltje and Perepelitsky in [BoPe20] introduced the notion of a Brauer pair of a $p$-permutation equivalence $\gamma$ (to be precise, they initially defined the notion for an element $\gamma\in T^{\Delta}(A,B)$ for which $\gamma\cdot_{B}\gamma^{*}=[A]$).
Because a $p$-permutation equivalence can be written as a difference of bimodules whose indecomposable direct summands have twisted diagonal vertices, applying the Brauer construction at a subgroup $X\leq G\times H$ to a $p$-permutation equivalence will only be nonzero when $X$ is twisted diagonal. However, this is not a sufficient criteria for the result to be nonzero, and this motivates the definitions given in this section. We review the definitions and results in [BoPe20].
@@
[BoPe20, Definition 5.1] Let $M\in{}_{kG}\textbf{triv}$. We call a Brauer pair $(P,e)$ of $kG$ an $M$-Brauer pair if $M(P,e):=e\cdot M(P)\neq 0$, where $M$ has $k[N_{G}(P,e)]e$-module structure. We denote the set of Brauer pairs of $M$ by $\mathcal{BP}(M)$.
@@
[BoPe20, Proposition 5.3, 5.4] Let $M,N\in{}_{kG}\mathbf{triv}$.
@@
    Let $M$ and $N$ be indecomposable $p$-permutation kGkG-modules, suppose that $(P,e)\in\mathcal{BP}(kG)$ is both a maximal $M$-Brauer pair and a maximal $N$-Brauer pair, and set $I:=N_{G}(P,e)$. Then $M(P,e)$ and $N(P,e)$ are indecomposable $p$-permutation kIekIe-modules. Moreover, $M\cong N$ if and only if $M(P,e)\cong N(P,e)$ as $kIe$-modules.
@@
[BoPe20, Definition 9.6] Let $\omega\in T(kG)$. A Brauer pair $(P,e)$ of $kG$ is called an $\omega$-Brauer pair if $\omega(P,e)=e\omega(P)\neq 0$ in $T(k[N_{G}(P,e)])$. The set of $\omega$-Brauer pairs is denoted $\mathcal{BP}(\omega)$.
@@
    If $\gamma\in T^{\Delta}(A,B)$ is a $p$-permutation equivalence for $A$ and $B$ and $\Delta(P,\phi,Q)\subset G\times H$ is a twisted diagonal subgroup, then via restriction, one can also view $\gamma(\Delta(P,\phi,Q))\in T(k[N_{G\times H}(\Delta(P,\phi,Q))])$ as an element of the corresponding Grothendieck group of $C_{G}(P)\times C_{H}(Q)=C_{G\times H}(\Delta(P,\phi,Q))$.
  2. (b)
@@
The following proposition gives convenient reformulations of being a Brauer pair for a $p$-permutation equivalence.
@@
[BoPe20, Proposition 10.8] Suppose $\gamma$ is a $p$-permutation equivalence for $A$ and $B$. Let $(\Delta(P,\phi,Q),e\otimes f^{*})\in\mathcal{BP}^{\Delta}(A,B)$ and set $I:=N_{G}(P,e)$ and $J:=N_{H}(Q,f)$. The following are equivalent:
@@
[BoPe20, Corollary 10.9] Let $(P,e)$ be an $A$-Brauer pair and define $\Lambda$ as in Theorem 6.7 and let $\Lambda_{B}\subseteq\Lambda$ be the set of pairs $(\psi,(Q,f))$ such that $(Q,f)$ is a $B$-Brauer pair. There exists a unique $H$-orbit of pairs $(\psi,(Q,f))\in\Lambda_{B}$ with the property that $(\Delta(P,\psi,Q),e\otimes f^{*})$ is a $\gamma$-Brauer pair.
@@
[BoPe20, Theorem 10.10] Suppose $\gamma$ is a $p$-permutation equivalence for $A$ and $B$. Let $\mathcal{I}$ denote the set of primitive idempotents in $Z(A)$ and let $\mathcal{J}$ denote the set of primitive idempotents in $Z(B)$.
@@
For claims (a) and (b), see [BoPe20, Theorem 10.10]. For (c), let $\lambda:=(\Delta(P,\phi,Q),b\otimes c^{*})$ be a Brauer pair for $k[G\times H]$. Then
@@
[BoPe20, Proposition 11.1] Let $(\Delta(P,\phi,Q),e\otimes f^{*})$ be a $\gamma$-Brauer pair and set $I:=N_{G}(P,e)$ and $J:=N_{H}(Q,f)$. For every $g\in I$ there exists a unique element $hC_{H}(Q)\in J/C_{H}(Q)$ such that $c_{g}\circ\phi=\phi\circ c_{h}:Q\to P$. Similarly for every $h\in J$, there exists a unique element $gC_{G}(P)\in I/C_{G}(P)$ such that $c_{g}\circ\phi=\phi\circ c_{h}$. These associations define mutually inverse group isomorphisms between $I/C_{G}(P)$ and $J/C_{H}(Q)$. The isomorphism $J/C_{H}(Q)\cong I/C_{G}(P)$ restricts to an isomorphism $QC_{H}(Q)/Q\cong PC_{G}(P)/P,hC_{H}(Q)\mapsto\phi(h)C_{G}(P)$, for $h\in Q$. The group $Y:=N_{G\times H}(\Delta(P,\phi,Q),e\otimes f^{*})$ satisfies $p_{1}(Y)=I,p_{2}(Y)=J,k_{1}(Y)=C_{G}(P),k_{2}(Y)=C_{H}(Q)$.
@@
[BoPe20, Theorem 11.4] Let $(\Delta(P,\phi,Q),e\otimes f^{*})\in\mathcal{BP}(\gamma)$ be a $\gamma$-Brauer pair and set $I:=N_{G}(P,e)$, $J:=N_{H}(Q,f)$. Suppose $C_{G}(P)\leq S\leq I$ and $C_{H}(Q)\leq T\leq J$ are intermediate groups related via the isomorphism in the previous proposition, and set $Y:=N_{S\times H}(\Delta(P,\phi,Q),e\otimes f^{*})=N_{S\times T}(\Delta(P,\phi,Q))$. Then the element $e\gamma(P,\phi,Q)f\in T(kY(e\otimes f^{*}))$ after restriction to $Y$ satisfies
@@
[BoPe20, Proposition 11.5] Let $(\Delta(D,\phi,E),e_{D}\otimes f^{*}_{E})$ be a maximal $\gamma$-Brauer pair. In particular, $D$ is a defect group of $A$ and $E$ is a defect group of $B$. For any subgroup $P\leq D$, let $(P,e_{P})$ denote the unique $A$-Brauer pair with $(P,e_{P})\leq(D,e_{D})$ and for any subgroup $Q\leq E$ let $(Q,f_{Q})$ denote the unique $B$-Brauer pair with $(Q,f_{Q})\leq(E,f_{E})$. Let $Q\leq E$ and set $P:=\phi(Q)$. Then $(\Delta(P,\phi,Q),e_{P}\otimes f^{*}_{Q})$ is a $\gamma$-Brauer pair.
@@
8. Brauer pairs for splendid Rickard equivalences
Each splendid Rickard equivalence $C$ induces a $p$-permutation equivalence $\gamma:=\Lambda(C)$, and therefore has an associated set of $\gamma$-Brauer pairs. As with $p$-permutation equivalences, in general applying the Brauer construction at an arbitrary subgroup to $C$ will not yield a local splendid Rickard equivalence. We may define a Brauer pair for a splendid Rickard equivalence in the same way as we have for $p$-permutation equivalences, treating contractible complexes as zero objects. In fact, the Lefschetz invariant of a contractible complex is zero in the corresponding Grothendieck group, but the converse of the statement rarely holds.
We will show that applying the Brauer construction at the Brauer pairs for the corresponding $p$-permutation equivalence $\gamma$ do indeed yield local splendid equivalences. Moreover, we demonstrate that applying the Brauer construction at any non-Brauer pair results in a contractible complex. The main result of this paper is the corollary of the two above results, which is that the Brauer pairs for a splendid Rickard equivalence and its corresponding $p$-permutation equivalence are the same.
For this section, suppose $A,B$ are direct summands of group algebras $kG$ and $kH$ respectively. Let $C$ be a splendid Rickard equivalence inducing a splendid equivalence between $A$ and $B$, and set $\gamma:=\Lambda(C)$. Given a Brauer pair $\omega=(\Delta(P,\phi,Q),e\otimes f^{*})$, denote by $\omega^{\circ}$ the “opposite” Brauer pair $(\Delta(Q,\phi^{-1},P),f\otimes e^{*}).$
@@ 8. Brauer pairs for splendid Rickard equivalences
Let $C$ be a splendid Rickard equivalence. The goal of this section is to determine the set $\mathcal{B}\mathcal{P}(C)$ of $C$-Brauer pairs. Notice that the Lefschetz invariant of $C(\omega)$ is $\gamma(\omega)$. Since the Lefschetz invariant of a contractible complex is zero, it is immediate that $\mathcal{BP}(\gamma)\subseteq\mathcal{BP}(C)$. However, a noncontractible complex can have Lefschetz invariant equal to zero, so it is not clear whether or not these sets are equal.
@@
First, we wish to reduce to the case that $A$ and $B$ are blocks. In order to do this, we will establish an analogue of Proposition 7.7 for splendid Rickard equivalences. The main tool is the following result.
@@
Since $C$ is a splendid Rickard equivalence, we have a homotopy equivalence $C\otimes_{B}C^{*}\simeq A$ of complexes of $(A,A)$-bimodules. We apply the functor $(\Delta(P),e\otimes e^{*})$ to this equation. Using Proposition 9.3, Theorem 6.7(d), and Proposition 3.4, we obtain an isomorphism
@@
Suppose $A$ and $B$ are sums of blocks and that $C$ is a splendid Rickard equivalence for $A$ and $B$. Let $f:\mathcal{I}\to\mathcal{J}$ be the bijection determined by $\gamma:=\Lambda(C)$ from Proposition 7.7.
@@
    We have $C=\bigoplus_{e\in\mathcal{I}}eCf_{e}$ in $K^{b}({}_{A}\mathbf{mod}_{B})$. Moreover, for every $e\in\mathcal{I}$, $eCf_{e}$ is a splendid Rickard equivalence for $kGe$ and $kGf_{e}$.
@@
The Lefchetz invariant of $eCf_{e}$ is $e\gamma f_{e}$, which is nonzero by Proposition 7.7(a). Therefore, $eCf_{e}\not\simeq 0$. By applying Lemma 8.3 to the splendid Rickard equivalence $C$ and the $A$-Brauer pair $(1,e)$, we obtain that there exists a unique $f\in\mathcal{J}$ such that $eCf\not\simeq 0$. By uniqueness, $f=f_{e}$. This proves (a).
@@
Therefore, it suffices to consider only splendid Rickard equivalences for block algebras. We remark that part (a) similarly holds for derived Rickard equivalences and stable equivalences of Morita type. Indeed, one may verify that a chain complex of bimodules inducing a derived equivalence between sums of block algebras decomposes into a direct sum of indecomposable derived Rickard complexes inducing equivalences of block algebras, and similarly for stable equivalences of Morita type.
From here, assume $A$ and $B$ are block algebras. We next prove a chain complex-theoretic version of Theorem 7.10.
@@
Suppose that $A$ and $B$ are blocks and $C$ is a splendid Rickard equivalence for $A$ and $B$, and set $\gamma:=\Lambda(C)$. Set $\lambda:=(\Delta(P,\phi,Q),e\otimes f^{*})$ be a $\gamma$-Brauer pair and set $I:=N_{G}(P,e),J:=N_{H}(Q,f)$. Suppose $C_{G}(P)\leq S\leq I$ and $C_{H}(Q)\leq T\leq J$ are intermediate groups related via the isomorphism $I/C_{G}(P)\to J/C_{H}(Q)$ defined in Proposition 7.9, and set $Y:=N_{S\times H}(\Delta(P,\phi,Q),e\otimes f^{*})$. Then the chain complex $C(\lambda):=eC(\Delta(P,\phi,Q))f\in\mathrm{Ch}^{b}({}_{kY(e\otimes f^{*})}\mathbf{triv})$ satisfies
@@
In particular, $C(\lambda)$ is a splendid Rickard equivalence for $k[C_{G}(P)]e$ and $k[C_{H}(Q)]f$, and the chain complex
@@
is a splendid Rickard equivalence for $kSe$ and $kTf$.
@@
of chain complexes of $k[N_{S\times S}(\Delta(P))](e\otimes e^{*})$-modules. Notice that when $\tau=(\phi,(Q,f))$, we have $X(\tau)=Y$, $C(\tau)=C(\lambda)$, and, by Proposition 7.9, $I(\tau):=N_{(S,\phi,J)}=S$. It follows from the latter that
@@
Thus, $C(\lambda)\otimes_{H}^{Y,Y^{\circ}}C(\lambda)^{*}$ is a noncontractible direct summand of the complex on the right hand side. On the other hand, $k[C_{G}(P)]e$ is an indecomposable $k[N_{S\times S}(\Delta(P))]$-bimodule, hence indecomposable and noncontractible as a chain complex. By the Krull-Schmidt theorem, precisely one of the terms in the direct sum must be noncontractible. We conclude that $k[C_{G}(P)]e\simeq C(\lambda)\otimes_{H}^{Y,Y^{\circ}}C(\lambda)^{*}$.
Therefore,
@@
The second homotopy equivalence follows similarly, and therefore $\operatorname{Ind}^{S\times T}_{Y}(C(\lambda))$ is a splendid Rickard equivalence for $kSe$ and $kTf$. ∎
We next show that in addition, if $\omega:=(\Delta(P,\phi,Q),e\otimes f^{*})$ is not a $\gamma$-Brauer pair, then $C(\omega)\simeq 0$ as a complex of $k[N_{G\times H}(\omega)]$-modules. First, we make some easy observations. These are well-known, but we include proofs for convenience.
@@
[…]
@@
Suppose $C$ is a splendid Rickard equivalence for $A$ and $B$, with $\gamma=\Lambda(C)$. If $\omega:=(\Delta(R,\psi,Q),d\otimes f^{*})\in\mathcal{BP}^{\Delta}(A,B)$ is not a $\gamma$-Brauer pair, then $C(\omega):=dC(\Delta(R,\psi,Q))f$ is contractible as a chain complex of $k[N_{G\times H}(\Delta(R,\psi,Q),d\otimes f^{*})](d\otimes f^{*})$-modules. In particular, $\omega$ is not a $C$-Brauer pair.
@@
There exists an $A$-Brauer pair $(P,e)$ and an isomorphism $\phi:Q\to P$ such that $\lambda:=(\Delta(P,\phi,Q),e\otimes f^{*})$ is a $\gamma$-Brauer pair. Set $I=N_{G}(P,e),\ J=N_{H}(Q,f)$, $K=N_{G}(R,d)$, $X=N_{G\times H}(\lambda)$, and $Y=N_{G\times H}(\omega)^{\circ}$. By Theorem 8.5, $\operatorname{Ind}_{X}^{I\times J}(C(\lambda))$ is a splendid Rickard equivalence for $kIe$ and $kJf$. In order to prove the claim, it suffices to show that
@@
[…]